What is the heat Capacity of 375 ml of water
1
Expert's answer
2019-08-08T10:54:24-0400

The specific heat capacity of water

"c=4180\:\rm{J/(kg\cdot K)}"

So, the heat capacity of 375 ml of water

"C=cm=4180\:\rm{J/(kg\cdot K)}\times 0.375\:\rm{kg}"

"=1567.5\:\rm{J/ K}"